### v2

@@ 5 On the semismooth property of solution maps to inclusions
Thus, by [35, Theorem 9.2], we obtain that $\sigma$ is Lipschitz continuous on $U$ with constant $\gamma$. Hence, by virtue of (5.14) we have
$D^{*}\sigma(x)(y^{*})\subset\Sigma(x,\sigma(x),0)\ \forall x\in U\ \forall y^{*}\in\mathbb{R}^{m}.$ (5.17)
@@
$M:=\sup\{\|(g_{x}^{T},g_{y}^{T})\|\,\big{|}\,(g_{x}^{T},g_{y}^{T})\in\Phi(u,\sigma(u)),u\in\tilde{U}\}<\infty.$
@@
The implication (i)$\Rightarrow$(ii) follows immediately from Proposition 5.6. We prove the implication (ii)$\Rightarrow$(iii). Assuming that (iii) does not hold there are sequences $(z_{k}^{*},x_{k}^{*},y_{k}^{*})\in L_{k}^{*}\subset(\bigcup{\mathcal{S}}^{*}F)(x,y,z)$ with $\|y_{k}^{*}\|\leq 1$ for all $k$ and $\|(z_{k}^{*},x_{k}^{*})\|\to\infty$ as $k\to\infty$. Since $L^{*}_{k}$ is a subspace, we have $(\tilde{z}^{*}_{k},\tilde{x}^{*}_{k},\tilde{y}^{*}_{k}):=(z_{k}^{*},x_{k}^{*},y_{k}^{*})/\|(z_{k}^{*},x_{k}^{*},y_{k}^{*})\|\in L_{k}^{*}$ for all $k$. Taking into account that the metric space ${\cal Z}_{m,n+m}$ is compact, by possibly passing to a subsequence we can assume that $L_{k}^{*}$ converges to some $L^{*}\in{\cal Z}_{m,n+m}$ and $(\tilde{z}^{*}_{k},\tilde{x}^{*}_{k},\tilde{y}^{*}_{k})$ converges to some $(\tilde{z}^{*},\tilde{x}^{*},\tilde{y}^{*})\in L^{*}$ with $\|(\tilde{z}^{*},\tilde{x}^{*},\tilde{y}^{*})\|=1$. Since ${\mathcal{S}}^{*}F(x,y,z)$ is closed in ${\cal Z}_{m,n+m}$, there holds $L^{*}\in{\mathcal{S}}^{*}F(x,y,z)$ and because of $\tilde{y}_{k}^{*}/\|(z_{k}^{*},x_{k}^{*},y_{k}^{*})\|\to 0$ we have $\tilde{y}^{*}=0$. Hence $(\tilde{z}^{*},\tilde{x}^{*},0)\in L^{*}\in{\mathcal{S}}^{*}F(x,y,z)$ and $\|(\tilde{z}^{*},\tilde{x}^{*})\|=1$ showing $L^{*}\not\in{\cal Z}_{m,n+m}^{\rm reg}$. This proves the implication (ii)$\Rightarrow$(iii). The implication (iii)$\Rightarrow$(i) holds trivially. Finally, formula (5.20) is a consequence of the definitions of ${\rm scd\,reg}F$, $(\bigcup{\mathcal{S}}^{*}F)$ and $\kappa(L^{*})$. ∎
@@
Due to the definition of the SCD property around a point and Proposition 5.8 we can find an open neighborhood $U^{\prime}\times V^{\prime}\times W^{\prime}$ of $(\bar{x},\bar{y},0)$ such that for every $(x,y,z)\in U^{\prime}\times V^{\prime}\times W^{\prime}$ the mapping $F$ has the SCD property around $(x,y,z)$ and ${\rm scd\,reg\,}F(x,y,z)<\infty$. Next consider the open neighborhood $U^{\prime\prime}:=\{x\in\Omega\cap U^{\prime}\,\big{|}\,\sigma(x)\in V^{\prime}\}$ and choose an open neighborhood $U$ of $\bar{x}$ satisfying $\mathrm{cl}\,U\subset U^{\prime\prime}$. This choice of $U$ ensures that ${\rm scd\,reg\,}F(x,\sigma(x),0)<\infty$, $x\in U$. We now show that $\Psi$ is locally bounded at $x\in U$. Indeed, we have $\gamma:={\rm scd\,reg\,}F(x,\sigma(x),0)<\infty$ and, due to the continuity of $\sigma$ and Proposition 5.8, we can find a neighborhood $\tilde{U}\subset\Omega$ of $x$ satisfying ${\rm scd\,reg\,}F(x^{\prime},\sigma(x^{\prime}),0)<\gamma+1$ $\forall x^{\prime}\in\tilde{U}$. Hence, $\|X^{*}\|<\gamma+1$ for every $X^{*}\in\Psi(x^{\prime})$ and our claim is established. In order to show the last assertion, consider $x\in U$ such that $F$ is SCD semismooth at $(x,\sigma(x),0)$, set $\gamma:={\rm scd\,reg\,}F(x,\sigma(x),0)$ and choose $\tilde{\delta}>0$ such that ${\rm scd\,reg\,}F(x^{\prime},\sigma(x^{\prime}),0)<\gamma+1$ $\forall x^{\prime}\in{\cal B}_{\tilde{\delta}}(x)$. By definition of the SCD semismooth property we can find for every $\epsilon>0$ some $\delta\in(0,\tilde{\delta})$ such that for every $x^{\prime}\in{\cal B}_{\delta}(x)$, every subspace $L^{*}\in{\mathcal{S}}^{*}F(x^{\prime},\sigma(x^{\prime}),0)$ and every $p\in\mathbb{R}^{m}$ we have
@@
[…]
@@
$\mathrm{gph}\,\partial^{c}_{y}f(x,\cdot)$ =$=\{(y,y-x)\,\big{|}\,y\not\in[x^{-},x^{+}]\}\cup\{(y,-y)\,\big{|}\,y\in(x^{-},x^{+})\}$
($\qquad\cup(\{x^{-}\}\times[x^{-}-x,-x^{-}])\cup(\{x^{+}\}\times(-x^{+},x^{+}-x))$
@@
by means of the MATLAB routine fmincon, which was able to compute in every case the correct global solution. Since ||$\cdot$|| ⋅ | is G-semismooth, i.e., |$\partial$|$\cdot$|∂ | ⋅ |-ss, and $\sigma$ is $\Psi$-ss with $\Psi$ given by (5.22), these functions are also semismooth with respect to the selections
@@
of |$\partial$|$\cdot$|∂ | ⋅ | and $\Psi$, respectiveley. By means of Lemma 3.2 and Corollary 5.10 we obtain that $\theta$ is semismooth with respect to the mapping

### v3

@@ 5 On the semismooth property of solution maps to inclusions
Thus, by [36, Theorem 9.2], we obtain that $\sigma$ is Lipschitz continuous on $U$ with constant $\gamma$. Hence, by virtue of (5.16) we have
$\emptyset\not=D^{*}\sigma(x)(y^{*})\subset\Sigma(x,\sigma(x),0)\ \forall x\in U\ \forall y^{*}\in\mathbb{R}^{m}.$ (5.19)
@@
$M:=\sup\{\|(g_{x}^{T},g_{y}^{T})\|\,\big|\,(g_{x}^{T},g_{y}^{T})\in\Phi(u,\sigma(u)),u\in\tilde{U}\}$$\infty$.
@@
[…]
@@
[…]
@@
Note that the respective solution mapping $S(x)=\{y\,\big|\,0\in F(x,y)\}$ amounts to the C-stationary-point mapping of the nonsmooth program $\min_{y}f(x,y)$ parameterized by $x$. For every $x\in\mathbb{R}$ we have
@@
$\mathrm{gph}\,\partial^{c}_{y}f(x,\cdot)$ $=\{(y,y-x)\,\big|\,y\not\in[x^{-},x^{+}]\}\cup\{(y,-y)\,\big|\,y\in(x^{-},x^{+})\}$
$\qquad\cup(\{x^{-}\}\times[x^{-}-x,-x^{-}])\cup(\{x^{+}\}\times(-x^{+},x^{+}-x))$
@@
by means of the MATLAB routine fmincon, which was able to compute in every case the correct global solution. Since $|\cdot|$ is G-semismooth, i.e., $\partial|\cdot|$-ss, and $\sigma$ is $\Psi$-ss with $\Psi$ given by (5.24), these functions are also semismooth with respect to the selections
@@
of $\partial|\cdot|$ and $\Psi$, respectiveley. By means of Lemma 3.2 and Corollary 5.10 we obtain that $\theta$ is semismooth with respect to the mapping